hi
can you please help me with math
how do you solve 57=a-40
and tell me how you got there with the awnser

Answer by stanbon(48545)

solve 57=a-40
Add 40 to both sides of the equation to get:
57+40 = a-40+40
97 = a
Switch sides to get:
a = 97
================
Cheers,
Stan H.